# Federico Echenique and Farzad Pourbabaee

### v2

@@ 5.1 Brunn-Minkowski Inequality
Given that Euclidean volume is translation invariant, we can shift the measured sets in the above inequality by $f_{i}$, and thus achieve the bound in (3.4), and thereby (3.3).∎
Proof of Theorem 2.
Since $f=\{f_{i}:i\in I\}$ is weakly Pareto optimal, then there is no allocation $g\in\mathcal{F}_{\bm{1}}$ such that $g_{i}\succ_{i}f_{i}$ for every $i\in I$. That is $\mathcal{F}_{\bm{1}}\cap\prod_{i\in I}\mathcal{U}_{i}^{(0)}(f_{i})=\emptyset$. That in turn means the normalized endowment vector $\omega=\bm{1}$ is disjoint from $\mathcal{V}^{(\varepsilon)}\coloneqq\sum_{i\in I}\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ for all $\varepsilon\geq 0$. Since preferences are convex the approximate upper contour sets are convex, so is their sum $\mathcal{V}^{(\varepsilon)}$. Therefore, by the hyperplane separation theorem and monotonicity of preferences there exists a nonzero vector $p\in\mathbb{R}^{d}_{+}$ such that $p\cdot v\geq p\cdot\bm{1}=\lVert p\rVert_{1}$ for all $v\in\mathcal{V}^{(0)}$. That is $\bm{1}\in H^{-}(p;\lVert p\rVert_{1})$ and $\mathcal{V}^{(0)}\subseteq H^{+}(p;\lVert p\rVert_{1})$.
Now suppose $v\in\mathcal{V}^{(\varepsilon)}$. Then, there are $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}$ for all $i\in I$, such that $v=\sum_{i\in I}g_{i}$ and $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ For each $i\in I$, it holds that $(1-\varepsilon)g_{i}\in\mathcal{U}_{i}^{(0)}$ and hence $(1-\varepsilon)\sum_{i\in I}g_{i}\in\mathcal{V}^{(0)}$. Consequently, one has $p\cdot(1-\varepsilon)v\geq\lVert p\rVert_{1}$. That, in turn, implies $v\in H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$, and thereby $\mathcal{V}^{(\varepsilon)}\subseteq H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$. As a result of this set inclusion, for an arbitrary $r>0$, one obtains that
@@
where the last inequality follows as before, because $\min_{p\neq\bm{0}}\lVert p\rVert_{1}/\lVert p\rVert_{2}=1$. Now set $A\coloneqq\mathcal{V}^{(\varepsilon)}\cap\mathbb{B}(\bm{1},r)$ and $B\coloneqq H^{-}(p\,;\,\lVert p\rVert_{1})\cap\mathbb{B}(\bm{1},r)$. By the above inequality one has $\mathsf{dist}(A,B)\geq\varepsilon$. Since $p$ is a nonzero vector in $\mathbb{R}_{+}^{d}$, the subset $B$ covers at least half of the volume of $\mathbb{B}(\bm{1},r)$. So it must be that $\mathsf{Vol}(A)\leq\mathsf{Vol}(B)$. Therefore, Lemma 1 implies that $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(\bm{1},r))\leq\mathrm{e}^{-\varepsilon^{2}d/8r^{2}}$, thus proving the volume bound (3.6) that is equivalent to (3.5).∎
Proof of Corollary 2.
Let $\beta\coloneqq\mathrm{CRU}(f)$. Since $f$ is not weakly Pareto optimal, then $\beta<1$. By Debreu (1951), one has $\beta\omega\in\partial\overline{\mathcal{V}^{(0)}}$. Thus, $\omega$ is a minimal element of the closure of $\beta^{-1}\mathcal{V}^{(0)}$. Observe that for every $i$, one has
@@
Therefore $\beta^{-1}\,\mathcal{V}^{(0)}=\mathcal{V}^{(1-\beta)}$, and $\omega$ becomes a minimal element of $\mathcal{V}^{(1-\beta)}$ as well. This fact, in turn, means that $f$ is $(1-\beta)$-Pareto optimal. Thus the corollary now follows from Theorem 2.∎
@@ 5.3 Concentration and Isoperimetric Inequalities
$\mu^{+}(A)\coloneqq\liminf_{\delta\to 0}\,\frac{\mu(A^{\delta})-\mu(A)}{\delta}\,.$ (5.5)
@@
$\mathcal{I}_{\mu}(a)\coloneqq\inf_{1/2<\mu(A)\leq 1-a}\mu^{+}(A)\,.$ (5.6)
In many environments, where the probability measure satisfy some mild regularity conditions, there exist universal lower bounds for the Isoperimetric function. One particular case that is of interest to us is the following lemma.
@@
Let $u$ be the uniform measure on the probability simplex $\Delta_{d}$. That is ${u}(A)=\mathsf{Area}(A)/\mathsf{Area}(\Delta_{d})$, for every $A\subseteq\Delta_{d}$. Then, there exists a universal constant $c>0$ such that for $a\in[0,1/2)$:
@@
In the following, we use $u(\cdot)$ to refer to the uniform measure on $\Delta_{d}$. As a corollary of the previous lemma we show that if a subset covers at least half of the measure on $\Delta_{d}$, then the measure of its $\delta$-extension is very close to $1$.
@@
$u(A^{\delta})$ $\geq u(A)+\int_{0}^{\delta}u(A^{t})\,\mathrm{d}t$
@@
where the third inequality follows from (5.7). Define $z(0)\coloneqq u(A)$, and let $z:[0,\delta]\to\mathbb{R}$ be the solution to the following integral equation:
@@
An important consequence of this result, that lies at the core of the proof of Theorem 3, is that if two subsets in $\Delta_{d}$ have positive distance from each other, then the area of at least one of them must be exponentially smaller than $\mathsf{Area}(\Delta_{d})$. Intuitively, this resembles the separation argument in Lemma 1, although its proof is not a direct consequence of the BN inequality, and follows from the more elaborate construct of the aforementioned Isoperimetric lower bound.
@@
5.4 Proofs of the results in Section 3.3
Proposition 1 implies that if an allocation is $\varepsilon$-Pareto dominated, then the extension of the subjective belief sets have empty intersection. Hence, any arbitrary split of the agents’ index set $I$ into two groups results into two subsets whose extensions have empty intersection as well. Thus, we can employ Lemma 3 to conclude that at least one of these subsets should have a small volume.
Proof of Proposition 1.
@@
where the last inequality follows because $\tilde{\mu}_{i}\cdot\big{[}g_{i}(1-\varepsilon)-f_{i}\big{]}>0$. Summing over all $i$’s leads to
$\eta\cdot\sum_{i=1}^{n}\big{[}(1-\varepsilon)g_{i}-f_{i}\big{]}>-\delta\sum_{i=1}^{n}\lVert(1-\varepsilon)g_{i}-f_{i}\rVert\geq-\delta\bar{\omega}\rho=-\varepsilon\bar{\omega}\,.$
@@
By Proposition 1, if $f$ is $\varepsilon$-Pareto dominated, then $\bigcap_{i\in I}\mathcal{B}_{i}(f_{i})^{\delta}=\emptyset$. For an arbitrary $J\subset I$, one can readily verify that $\mathcal{B}_{J}^{\delta}\subseteq\bigcap_{j\in J}\mathcal{B}_{j}(f_{j})^{\delta}$. Therefore, $\mathcal{B}_{J}^{\delta}\cap\mathcal{B}_{J^{c}}^{\delta}=\emptyset$. Applying Lemma 3 implies the conclusion in (3.8).∎
Appendix A Remaining Proofs
Continuity of the Approximate Pareto Correspondence
Fix an exchange economy $\mathcal{E}$ and denote the set of all $\varepsilon$-Pareto optimal allocations by $\bm{P}^{(\varepsilon)}$. The subset $\bm{P}^{(\varepsilon)}$ is evidently increasing with respect to the set inclusion order: $\bm{P}^{(\varepsilon^{\prime})}\subseteq\bm{P}^{(\varepsilon^{\prime\prime})}$ for $\varepsilon^{\prime}\leq\varepsilon^{\prime\prime}$. Using continuity and monotonicity of preferences, we show that the correspondence $\bm{P}:[0,1]\twoheadrightarrow\mathbb{R}_{+}^{d\times I}$ is upper-hemicontinuous on $[0,1]$ and continuous at $\varepsilon=0$. This means as $\varepsilon\to 0$ the subset $\bm{P}^{(\varepsilon)}$approximates” the space of weakly Pareto optimal allocations.
Proposition A.1.
The correspondence $\bm{P}:[0,1]\twoheadrightarrow\mathbb{R}_{+}^{d\times I}$ is upper-hemicontinuous.
Proof.
Let $\varepsilon_{k}\to\varepsilon$, $f^{(k)}\in\bm{P}^{(\varepsilon_{k})}$. Since $\mathcal{F}_{w}$ is a compact subset of $\mathbb{R}_{+}^{d\times I}$, then $f^{(k)}$ has a limit point, i.e., $\exists f\in\mathcal{F}_{w}$ such that the subsequence $f^{(k_{m})}\to f$ as $m\to\infty$. To avoid clutter, we denote the indices of this subsequence by $k$ instead of $k_{m}$. To justify the upper-hemicontinuity, we need to show that $f\in\bm{P}^{(\varepsilon)}$. Since $f^{(k)}\to f$, then $\sum_{i\in I}^{n}\lVert f^{(k)}_{i}-f_{i}\rVert\to 0$, that in turn implies $f^{(k)}_{i}\to f_{i}$ uniformly over $i\in I$. Now assume by contradiction that $f\notin\bm{P}^{(\varepsilon)}$. Then, there exists $g\in\mathcal{F}_{w}$ such that $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ for all $i\in I$. Since preferences are continuous (A2), there exists a $\bar{k}$ such that $(1-\varepsilon)g_{i}\succ_{i}f_{i}^{(k)}$ for all $k\geq\bar{k}$ and for all $i\in I$. Because of continuity again, for each $i$ there exists $\hat{\varepsilon}_{i}>0$, such that $(1-\varepsilon-\hat{\varepsilon}_{i})g_{i}\succ_{i}f_{i}^{(k)}$ for all $k\geq\bar{k}$. Let $\hat{\varepsilon}\coloneqq\min_{i\in I}\hat{\varepsilon}_{i}$, then because of monotonicity (A3) one has $(1-\varepsilon-\hat{\varepsilon})g_{i}\succ_{i}f_{i}^{(k)}$ for all $k\geq\bar{k}$. Pick $k_{0}=\min\{k\geq\bar{k}:\varepsilon_{k}\leq\varepsilon+\hat{\varepsilon}\}$. Since $\varepsilon_{k}\to\varepsilon$, then $k_{0}$ is finite. Observe that the previous statement implies that $f^{(k_{0})}$ is not $\varepsilon_{k_{0}}$-Pareto optimal. Thus, the contradiction is reached and hence $f$ must belong to $\bm{P}^{(\varepsilon)}$. ∎
Corollary A.1 (Continuity at $\varepsilon=0$).
Suppose $\varepsilon_{k}\to 0$ and $f^{(k)}\in\bm{P}^{(\varepsilon_{k})}$. Then, there exists a subsequence $\{f^{(k_{m})}\}$ that converges to a weakly Pareto optimal allocation $f$. This is resulted from the upper hemicontinuity of $\bm{P}^{(\cdot)}$ at $\varepsilon=0$. Conversely, the correspondence is clearly lower-hemicontinuous at $\varepsilon=0$, because every weakly Pareto optimal allocation is in fact $\varepsilon$-Pareto optimal.
Proof of Proposition 2

### v4

@@ 5.1 Brunn-Minkowski Inequality
Since the Euclidean volume is translation invariant, we can shift the subsets in the above inequality by $f_{i}$, and thus obtain:
Subsequently, a measure change from the uniform law $\mathsf{P}^{r}$ to $\mathsf{P}^{r}_{\kappa}$ implies (3.3).∎
Proof of Theorem 2
Since $f=(f_{i})_{i\in I}$ is weakly Pareto optimal, then there is no allocation $g\in\mathcal{F}_{\bm{1}}$ such that $g_{i}\succ_{i}f_{i}$ for every $i\in I$. That is $\mathcal{F}_{\bm{1}}\cap\prod_{i\in I}\mathcal{U}_{i}^{(0)}(f_{i})=\emptyset$. That in turn means the normalized endowment vector $\omega=\bm{1}$ is disjoint from $\mathcal{V}^{(\varepsilon)}\coloneq\sum_{i\in I}\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ for all $\varepsilon\geq 0$. Since preferences are convex the approximate upper contour sets are convex, so is their sum $\mathcal{V}^{(\varepsilon)}$. Therefore, by the hyperplane separation theorem and monotonicity of preferences there exists a nonzero vector $p\in\mathbb{R}^{d}_{+}$ such that $p\cdot v\geq p\cdot\bm{1}=\lVert p\rVert_{1}$ for all $v\in\mathcal{V}^{(0)}$. That is $\bm{1}\in H^{-}(p;\lVert p\rVert_{1})$ and $\mathcal{V}^{(0)}\subseteq H^{+}(p;\lVert p\rVert_{1})$.
Now suppose $v\in\mathcal{V}^{(\varepsilon)}$. Then, there are $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}$ for every $i\in I$, such that $v=\sum_{i\in I}g_{i}$ and $(1-\varepsilon)g_{i}\succ_{i}f_{i}$ For each $i\in I$, it holds that $(1-\varepsilon)g_{i}\in\mathcal{U}_{i}^{(0)}$ and hence $(1-\varepsilon)\sum_{i\in I}g_{i}\in\mathcal{V}^{(0)}$. Consequently, one has $p\cdot(1-\varepsilon)v\geq\lVert p\rVert_{1}$. That in turn implies $v\in H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$, and thereby $\mathcal{V}^{(\varepsilon)}\subseteq H^{+}\left(p;\lVert p\rVert_{1}/(1-\varepsilon)\right)$. As a result of this set inclusion, for an arbitrary $r>0$, one obtains that
@@
where the last inequality follows as before, because $\min_{p\neq\bm{0}}\lVert p\rVert_{1}/\lVert p\rVert_{2}=1$. Now set $A\coloneq\mathcal{V}^{(\varepsilon)}\cap\mathbb{B}(\bm{1},r)$ and $B\coloneq H^{-}(p\,;\,\lVert p\rVert_{1})\cap\mathbb{B}(\bm{1},r)$. By the above inequality one has $\mathsf{dist}(A,B)\geq\varepsilon$. Since $p$ is a nonzero vector in $\mathbb{R}_{+}^{d}$, the subset $B$ covers at least half of the volume of $\mathbb{B}(\bm{1},r)$. So it must be that $\mathsf{Vol}(A)\leq\mathsf{Vol}(B)$. Therefore, Lemma 1 implies that $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(\bm{1},r))\leq\mathrm{e}^{-\varepsilon^{2}d/8r^{2}}$, thus proving the volume bound in (3.5) that is equivalent to (3.4) after a change of measure from $\mathsf{P}^{r}$ to $\mathsf{P}^{r}_{\kappa}$.∎
Proof of Proposition 1
Let $\beta\coloneq\mathrm{CRU}(f)$. Since $f$ is not weakly Pareto optimal, then $\beta<1$. By Debreu (1951), it holds that $\beta\bm{1}\in\partial\overline{\mathcal{V}^{(0)}}$. Observe that for every $i$, one has
@@
Therefore $\beta^{-1}\,\mathcal{V}^{(0)}(f)=\mathcal{V}^{(1-\beta)}(f)$, and $\bm{1}$ becomes a minimal element of $\mathcal{V}^{(1-\beta)}(f)$. Following the same recipe as in the proof of Theorem 2, let $p$ be the supporting vector of $\mathcal{V}^{(1-\beta)}$ at $\bm{1}$, and consider the $r$-ball centered at $\beta\bm{1}$. Set $A\coloneq\mathcal{V}^{(1-\beta)}\cap\mathbb{B}(\beta\bm{1},r)\subseteq H^{+}\left(p;\lVert p\rVert_{1}\right)$ and $B\coloneq H^{-}\left(p;\beta\lVert p\rVert_{1}\right)\cap\mathbb{B}(\beta\bm{1},r)\subseteq H^{-}\left(p;\beta\lVert p\rVert_{1}\right)$. Then, $\mathsf{dist}(A,B)\geq 1-\beta$, and $\mathsf{Vol}(B)\geq\mathsf{Vol}(A)$. Therefore by Lemma 1, $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(\beta\bm{1},r))\leq\mathrm{e}^{-(1-\beta)^{2}d/8r^{2}}$, that is equivalent to
$\mathsf{P}^{r}\left(\beta\bm{1}+z\in\mathcal{V}^{(1-\beta)}(f)\right)\leq\mathrm{e}^{-(1-\beta)^{2}d/8r^{2}}\,.$
Because $\beta^{-1}\mathcal{V}^{(1-\beta)}(f)=\mathcal{V}^{(1-\beta^{2})}(f)$, the above bound implies that
$\mathsf{P}^{r}\left(\bm{1}+\beta^{-1}z\in\mathcal{V}^{(1-\beta^{2})}(f)\right)\leq\mathrm{e}^{-(1-\beta)^{2}d/8r^{2}}\,.$
Since $z$ is uniform on the $r$-ball, then $\beta^{-1}z$ becomes uniform on the $\beta^{-1}r$-ball. Thus the above bound reduces to
$\mathsf{P}^{r}\left(\bm{1}+z\in\mathcal{V}^{(1-\beta^{2})}(f)\right)\leq\mathrm{e}^{-\left(\frac{1-\beta}{\beta}\right)^{2}d/8r^{2}}\,.$
By changing the measure from $\mathsf{P}^{r}$ to $\mathsf{P}^{r}_{\kappa}$ the inequality in (3.6) follows. ∎
@@ 5.3 Concentration and Isoperimetric Inequalities
$\mu^{+}(A)\coloneq\liminf_{\delta\to 0}\,\frac{\mu(A^{\delta})-\mu(A)}{\delta}\,.$ (5.6)
@@
$\mathcal{I}_{\mu}(a)\coloneq\inf_{1/2<\mu(A)\leq 1-a}\mu^{+}(A)\,.$ (5.7)
In many environments, where the probability measures satisfy some mild regularity conditions, there exist universal lower bounds for the isoperimetric function. One particular case that is of interest to us is the following lemma.
@@
Let $u$ be the uniform measure on the probability surface $\Delta_{d}$. That is ${u}(A)=\mathsf{Vol}(A)/\mathsf{Vol}(\Delta_{d})$, for every $A\subseteq\Delta_{d}$. Then, there exists a universal constant $c>0$ such that for $a\in[0,1/2)$:
@@
In the following, we use $u(\cdot)$ to refer to the uniform measure on $\Delta_{d}$. As a corollary of the previous lemma we show that if a subset covers at least half of the measure on $\Delta_{d}$, then the uniform measure of its $\delta$-extension is very close to $1$.
@@
$u(A^{\delta})$ $\geq u(A)+\int_{0}^{\delta}u^{+}(A^{t})\,\mathrm{d}t$
@@
where the third inequality follows from (5.8). Define $z(0)\coloneq u(A)$, and let $z:[0,\delta]\to\mathbb{R}$ be the solution to the following integral equation:
@@
An important consequence of this result, that lies at the core of the proof of Theorem 3, is that if two subsets in $\Delta_{d}$ have positive distance from each other, then the volume of at least one of them must be exponentially smaller than $\mathsf{Vol}(\Delta_{d})$. Intuitively, this resembles the separation argument in Lemma 1, although its proof is not a direct consequence of the Brunn-Minkowski inequality, and follows from the more elaborate construct of the aforementioned isoperimetric lower bound.
@@
5.4 Proof of the Results in Section 3.3
Proposition 2 implies that if an allocation is $\varepsilon$-Pareto dominated, then the extension of the subjective belief sets have an empty intersection. Hence, any arbitrary split of the agents’ index set $I$ into two groups results in two subsets whose extensions also have an empty intersection. Thus, we can employ Lemma 3 to conclude that at least one of these subsets should have a small volume.
Proof of Proposition 2
@@
where the last inequality follows because $\tilde{\mu}_{i}\cdot\big{[}g_{i}(1-\varepsilon)-f_{i}\big{]}>0$. Summing over all agents $i$ leads to
$\eta\cdot\sum_{i\in I}\big{[}(1-\varepsilon)g_{i}-f_{i}\big{]}>-\delta\sum_{i\in I}\lVert(1-\varepsilon)g_{i}-f_{i}\rVert\geq-\delta\bar{\omega}\rho=-\varepsilon\bar{\omega}\,.$
@@
By Proposition 2, if $f$ is $\varepsilon$-Pareto dominated, then $\bigcap_{i\in I}\mathcal{B}_{i}(f_{i})^{\delta}=\emptyset$, where $\delta=\varepsilon/\rho$. For an arbitrary $J\subset I$, one can readily verify that $\mathcal{B}_{J}^{\delta}\subseteq\bigcap_{j\in J}\mathcal{B}_{j}(f_{j})^{\delta}$. Therefore, $\mathcal{B}_{J}^{\delta}\cap\mathcal{B}_{J^{c}}^{\delta}=\emptyset$. Under the $\ell_{2}$-norm, equation (3.7) implies that $\rho=2\sqrt{d}$, and thus $\delta=\varepsilon/2\sqrt{d}$. Applying Lemma 3 together with the previous expression for $\delta$ lead to the conclusion in (3.8).∎
6 Remaining Proofs
6.1 Gaussian Measure on $\mathbb{B}(r)$
In Section 2.1 we remarked that the constant $\kappa$ for a conditional Gaussian measure is independent of $d$. Here we include a proof of this fact.
Let $\gamma_{d}\sim\mathcal{N}(0,\bm{I}_{d})$ be the standard $d$-dimensional Gaussian measure, and $\widetilde{\mathsf{P}}^{r}$ be its restriction to the Euclidean $r$-ball. Then the Radon-Nikodym derivative of $\widetilde{\mathsf{P}}^{r}$ with respect to the uniform measure $\mathsf{P}^{r}$ on the $r$-ball is
$\begin{gathered}\frac{\mathrm{d}\widetilde{\mathsf{P}}^{r}}{\mathrm{d}\mathsf{P}^{r}}(z)=\frac{\mathrm{d}\gamma_{d}(z)}{\mathrm{d}z}\,\,\frac{\int_{\mathbb{B}(r)}\mathrm{d}x}{\int_{\mathbb{B}(r)}\mathrm{d}\gamma_{d}(x)}=\mathrm{e}^{-\lVert z\rVert^{2}/2}\,\,\frac{\int_{\mathbb{B}(r)}\mathrm{d}x}{\int_{\mathbb{B}(r)}\mathrm{e}^{-\lVert x\rVert^{2}/2}\,\mathrm{d}x}\leq\frac{\int_{\mathbb{B}(r)}\mathrm{d}x}{\int_{\mathbb{B}(r)}\mathrm{e}^{-\lVert x\rVert^{2}/2}\,\mathrm{d}x}\,.\end{gathered}$
Let us denote the radial component of $x$ by $\rho$, i.e., $\rho=\lVert x\rVert$. Then, $0\leq\rho\leq r$ and one has
$\frac{\int_{\mathbb{B}(r)}\mathrm{d}x}{\int_{\mathbb{B}(r)}\mathrm{e}^{-\lVert x\rVert^{2}/2}\,\mathrm{d}x}=\frac{\int_{0}^{r}\rho^{d-1}\,\mathrm{d}\rho}{\int_{0}^{r}\mathrm{e}^{-\rho^{2}/2}\rho^{d-1}\,\mathrm{d}\rho}\leq\mathrm{e}^{r^{2}/2}\,,$
where in the last inequality, $\mathrm{e}^{-\rho^{2}/2}$ (in the integrand of the denominator) is lower bounded by $\mathrm{e}^{-r^{2}/2}$. Hence, the Radon-Nikodym derivative of the restricted standard Gaussian on the $r$-ball relative to the uniform measure is uniformly bounded above by $\mathrm{e}^{r^{2}/2}$.
6.2 Proof of Proposition 3